Prev Next

SW&IT Meeting Minutes 2024-06-24

Attendees: Brandon, Glenn, LeRoy, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • xxx
  • This week
    • xxx

YZ

  • Previous:
    • Patch and test MODS real time image verification service (new version deployed, unfortunately still has a bug).
    • MODS script engines patch tests with Olga (all good and worked as expected, IT#6939 updated).
    • TMS software refactoring (TMS laser module still en route to Tucson).
  • Next:
    • Patch and redeploy MODS image verification service again (needs IT help).
    • TMS software refactoring.
    • Investigate IT#4236 (LBC housekeeping hang).
    • Some discussions on MODS/LUCI ETC (and their implementations).
    • Planning on refactoring LBC software to left/right IIF proxies.
    • Other misc items (planning on more FPIA tests, NOMICS image packing, OSCO+/DISCO etc).

GCE

  • Last Week
    • New IT 9130. The pointing log stops when the instrument authorization is changed. Am looking into changing this behavior.
    • Finished investigation of the source of the pk libraries currently in use by the PCS on the TCS system. The pk libraries in tcs-test:/lbt/tcs_devel/pk-2.3.3 do, in fact match those contained in the pk GitHub repository, except for compilation options.
    • Successfully compiled GEIRS r806M-13. Will install next ECD.
    • Built a VM for work on LUCI ETC.
  • This Week
  • Carryover
    • Observer GUI: Fix “More Details” and “Edit Item” menus, which were never finished and do not work, and have been disabled for now.

PSG

  • Last week
    • Holiday
    • Had instance of AZCAM server crash.
      • Additional logging shows where it happened
      • Looks like software tried to talk to a disconnected camera.
      • Looking into a fix.
    • Fixed LBTI 'setstars' to remove guidestar (not 'live' yet).
    • SHARK discussion about work to be done during shutdown.
      • Fresh install of software on new HDD.
      • Camera feature mods (involves coordination with Jordan)
  • This week
    • Meeting for work on new FFTcam.
    • Continue working on AZCAM server crash.
    • Talk about new refmem entries for LBTI/SHARK OBJNAME.
    • LMIRCAM updates to be discussed.
    • Day off Tuesday?

LTD

  • Last week:
    • Configured tcs server replacements
  • This week:

BMM

  • last week:
    • ocam zero pixel: trigger reset_ocam
    • doug's AO spectroscopy using TT offset
    • ao web gui:
      • fix to processes backend to avoid conflicts with w_check
      • lots of small frontend fixes, refactoring
  • this week:
    • finish up ocam zero pixel / reset_ocam, test plan
    • select correct autogain.conf based on ACE/TTM mode
    • cacao benchmarks
    • ao web gui operate button / progress

SH

Topic revision: r7 - 24 Jun 2024, PaulGrenz
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback